Which is closer to 1/2: 8/15 or 3/7? Explain your reeasoning
Answer by Alan3354(69443)

Which is closer to 1/2: 8/15 or 3/7?
--------
Method 1
1/2 = 105/210
3/7 = 90/210
8/15 = 112/210
------------------
3/7 - 1/2 = -15/210
8/15 -1/2 = +7/210
////////////////////
Method 2
1/2 = 0.5
3/7 = 0.42857, diff = 0.07143
8/15 = 0.53333, diff= 0.03333